# CaVE: A Cone-Aligned Approach for Fast Predict-then-optimize with Binary Linear Programs

### 12 Dec 2023 v1

@@ 4.2 Cone-aligned Vector Estimation: three variants
$\mathcal{L}_{\text{CaVE}}(\hat{\bm{c}},\widetilde{\bm{A}})=-\text{cosine\_similarity}(\hat{\bm{c}},\bm{p}_{\hat{\bm{c}}})=-\dfrac{\hat{\bm{c}}\cdot\bm{p}_{\hat{\bm{c}}}}{\lVert\hat{\bm{c}}\rVert_{2}\cdot\lVert\bm{p}_{\hat{\bm{c}}}\rVert_{2}},$ (4)
where, assuming there are $m$ binding constraints, i.e., $\widetilde{\bm{A}}\in\mathbb{R}^{m\times d}$, the projection writes:
$\bm{p}_{\hat{\bm{c}}}=\bm{\lambda}^{*}\widetilde{\bm{A}},\quad\bm{\lambda}^{*}=\operatorname*{arg\,min}_{\bm{\lambda}\geq\bm{0}}\lVert\bm{\lambda}\widetilde{\bm{A}}-\hat{\bm{c}}\rVert^{2}.$ (5)
@@
6:        Compute loss $\mathcal{L}_{\text{CaVE}}(\hat{\bm{c}},\widetilde{\bm{A}})$ (4)
@@
Figure 3 illustrates the three types of projection that can be employed in the CaVE framework. The exact projection projects cost coefficients $\hat{\bm{c}}$ directly onto the surface of the optimal subcone; this is the approach that we have just laid out in Algorithm 1. The inner projection ensures that the predicted vector lies strictly within the subcone; and The heuristic projection is an approximation of the true projection onto the optimal subcone, used to reduce the computational cost. We will detail these three variants next.
@@
CaVE-E performs exact projection, wherein the optimal solution of quadratic programming problem (5) is computed to locate the projection on the cone’s surface, as illustrated on the left in Fig. 3. Nevertheless, the CaVE-E method encounters a significant drawback due to its projection onto the face of the cone and the use of cosine similarity as the loss function. This approach results in the vanishing of gradients as the predicted cost vector nears the surface of the optimal subcone but is yet to enter it. Experimental evidence in Section 6 corroborates this issue – the regret of CaVE-E is typically higher than existing end-to-end methods, necessitating a modification.
@@
Due to the issue of vanishing gradients associated with CaVE-E, CaVE+ replaces the exact projection with what we refer to as “inner projection”. The goal is to obtain a projection of the predicted cost vector that lies inside the subcone. As all optimal projections will lie on a face of the subcone, we thus require a suboptimal solution to the projection problem. This is readily achieved by simply limiting the number of iterations in the quadratic programming solver and thus terminating prematurely. In our experiments, the maximum iteration of CaVE+ was set to $3$. A suboptimal projection will lie inside the optimal subcone, resulting in non-zero loss and a strong gradient signal that will push the ML model parameters to produce predictions that move toward the inside of the subcone. Compared to exact methods, this approach also has computational efficiency. This approach is illustrated in the middle in Fig. 3.
@@
To alleviate the computational burden of repeated QP solving in both CaVE-E an CaVE+, a hybrid strategy is employed in CaVE-H. We interleave inner projections (obtained with a QP just as in CaVE+) with much cheaper heuristic projections. Unlike exact and inner projections, the heuristic projection does not necessitate solving a quadratic program and instead requires a simple convex combination of $\overline{\bm{A}}$ and $\hat{\bm{c}}$ with weight $\alpha\in(0,1)$:
$\bm{p}_{\hat{\bm{c}}}=(1-\alpha)\hat{\bm{c}}+\alpha\overline{\bm{A}},$
where $\overline{\bm{A}}=\dfrac{1}{m}\sum_{j=1}^{m}{\widetilde{\bm{A}}_{j}}$ is the average binding constraints. As illustrated on the right in Fig. 3, it is crucial to note that the heuristic projection is not guaranteed to be in the optimal subcone, but it still ensures that the cost coefficient vector is pushed in the direction of the optimal subcone. With probability $\beta<0.5$, i.e., in the minority of the iterations of Algorithm 1CaVE-H performs an inner projection via QP. With probability $(1-\beta)$, the heuristic projection is used instead, without any optimization required. In our experiments, $\alpha$ and $\beta$ are set to $0.2$ and $0.3$, respectively, and they were not tuned any further for performance.
SPO+ PFYL NCE CaVE-E CaVE+ CaVE-H
Cost per iteration 1×1$\times$1 ×BLP $K\times$BLP $\beta\times$BLP 1×1$\times$1 ×QP 1×1$\times$1 ×QP (partial) $\beta\times$QP (partial)
Table 1: Comparison of state-of-the-art predict-then-optimize methods SPO+, PFYL and NCE with the three CaVE variants w.r.t. “Cost per iteration” (per training instance), i.e., the frequency of time-consuming solver calls required to compute a method’s loss during gradient descent. SPO+ requires solving the BLP with the predicted costs, and PFYL requires solving the $K$ (typically 1-5) BLPs, whereas NCE needs to solve only a small fraction of the BLPs. CaVE methods require solving a single QP, partially solving a single QP, or partially solving a single QP $100\times\beta\%$ of the time.
@@
Table 1 summarizes the computational cost of training predict-then-optimize models using SPO+ [7], PFYL [3], NCE [20] and the three CaVE variants. We choose these SPO+ and PFYL as they represent the state-of-the-art based on an extensive evaluation carried out in [28], though our experiments will include NCE, known as a fast method, as well as accelerated variants of SPO+ and PFYL. The table shows that CaVE-H and NCE are the fastest to train whereas PFYL and SPO+ exhibit similar costs, particularly when PFYL’s number of random perturbations of the predicted cost vector is set to $K=1$CaVE+ sits between CaVE-E and CaVE-H in terms of training cost. We will empirically examine these theoretical complexities in the next section.
@@ 5 Benchmark Datasets
The shortest path problem is a fundamental task in graph theory and network analysis, where the objective is to find a minimum weight path from a fixed source node to a fixed target node. Our instances are based on $5\times 5$ grid networks (SP5) where the source is the node in the northwest corner of the grid and the target is the node in the southeast corner. The cost coefficient $c_{ij}$ come from
@@
$\Bigg{(}\frac{1}{\sqrt{10}}\left(\mathcal{B}\bm{x}_{i}\right)_{j}+3\Bigg{)}^{\text{deg}}\cdot\epsilon_{ij},$
@@ 6.1 Experimental setup
The experiments were designed to evaluate both the training time and normalized regret, with the size of both the training and test sets being 1,000 instances. To account for randomness in data generation and stochastic model training, $10$ random seeds are used to generate and train $10$ training/test sets and corresponding ML models. Our comparative analysis encompassed several methods, including three variants of CaVE, a two-stage approach (least-squares regression on cost vectors), SPO+, PFYL, and NCE. A linear model was used as the cost prediction model class for all the aforementioned approaches; this is a standard choice in the literature starting with [7].
The Adam optimizer was used for gradient descent for all our experiments. For SP5, a learning rate of $0.01$ was used with all methods for 10 epochs, except for the two-stage and NCE method which allotted 20 epochs at the same learning rate. For both TSP and CVRP, the only modification was to increase the learning rate to $0.05$, maintaining the number of epochs at 10. In selecting the number of epochs, careful consideration was given to the convergence of each model, ensuring that our results are not skewed by over or under-training. The loss curves in Fig. 4 show that all methods exhibit a convergent behaviour, with CaVE methods doing so earlier than the baselines. Note that the 2-Stage models are trained essentially to optimality as they are convex optimization problems and are thus omitted from Fig. 4.
Figure 4: Training loss curves for TSP50 (top) and CVRP20 (bottom) with a polynomial of degree 4. The vertical axis represents the loss values of each of the five methods, normalized to the $[0,1]$ range where 1 and 0 correspond to the maximum and minimum loss values encountered by each method for 320 mini-batch gradient descent iterations (10 epochs). Normalization is necessary because each method has its own loss function with values on different orders of magnitude.
@@ 6.2 Results
Tables 2(b)3(b)4(b)5(b), and 6(b) summarize the results. The tables correspond to the Shortest Path problem on $5\times 5$ grid (SP5), TSP with 20 nodes (TSP20), TSP with 50 nodes (TSP50), CVRP with 20 customers (CVRP20), and CVRP with 30 customers (CVRP30), respectively. Each table contains two sub-tables: (a) reports the normalized regret metric defined in Section 3.2 on 1,000 test instances and (b) reports the training time in seconds; both quantities are averages over $10$ random seeds. Each sub-table has two rows, one corresponding to datasets that use a degree-4 polynomial in the (unknown) mapping from instance features to cost coefficients, and the other for a degree-6 polynomial; we refer to Section 5 for details on the role of the polynomial in the function that we are attempting to learn, but note that the higher the degree the more difficult the learning task. In each sub-table, we bold the best-performing method, excluding the 2-Stage method as it is often fast to train but has much worse regret than end-to-end methods.
@@
This is the easiest BLP (in fact, LP) we will look at. It serves as a sanity check for any new method in this space. Table 2(b) shows that SPO+ and PFYL achieve the lowest test regrets, closely followed by CaVE-H and CaVE+. The latter two train roughly $10$ times faster than SPO+ and PFYL, already substantiating our claim that QP solving is faster in solving the optimization problem itself, even for small-scale polynomial-time solvable shortest path problems. In addition, although NCE is fast, it has higher regret than others.
@@
(a) Average Test Normalized Regret
@@
Deg $4$ 0.0882 0.1073 0.0839 0.0835 0.0779 0.0768 0.1134
Deg $6$ 0.1258 0.1130 0.0889 0.0884 0.0772 0.0786 0.1378
(b) Average Training Time (Sec)
@@
For the TSP, SPO+, PFYL and NCE employ the Dantzig–Fulkerson–Johnson (DFJ) formulation [5] to solve this BLP efficiently. For both TSP20 and TSP50, CaVE+ achieves the best time-regret tradeoff across all methods: its regret is the best or second-best across all methods and its training time is the second-best after CaVE-H. As shown in part (b) of Tables 3(b) and 4(b)CaVE+ trains in roughly half of the time of SPO+ and PFYL, achieving the same or better test regret. Additionally, we also compared our methods with SPO+ Rel and PFYL Rel, which employ a linear relaxation of the BLP during training, as detailed in the Appendix 0.A.1. While they do reduce the training time of vanilla SPO+PFYL by a bit, this typically comes at at an increase in regret. Similarly, NCE has a prohibitively high regret.
@@
(a) Average Test Normalized Regret
@@
Deg $4$ 0.1212 0.0735 0.0620 0.0769 0.0595 0.0656 0.1220
Deg $6$ 0.2132 0.0801 0.0697 0.0952 0.0748 0.0741 0.1485
(b) Average Training Time (Sec)
@@
(a) Average Test Normalized Regret
@@
Deg $4$ 0.2816 0.1519 0.0769 0.0959 0.0756 0.0803 0.1431
Deg $6$ 0.5261 0.2325 0.0857 0.1128 0.1026 0.0898 0.1709
(b) Average Training Time (Sec)
@@
(a) Average Test Normalized Regret
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL
Deg $4$ 0.1036 0.0917 0.0735 0.0798 0.0587 0.0708
Deg $6$ 0.1967 0.0924 0.0757 0.0967 0.0827 0.0886
(b) Average Training Time (Sec)
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL
Deg $4$ 1.76 189.95 147.20 46.32 4665.73 4705.95
Deg $6$ 1.92 196.98 172.78 50.64 4029.40 4991.41
As mentioned earlier, we are the first to tackle a CVRP in an end-to-end predict-then-optimize setting. To solve CVRP as a BLP, we formulated the problem with k-path cuts [12] and solved it using Gurobi. For CVRP20, CaVE+ and SPO+ compete for the lowest regret, with CaVE-H and PFYL running closely behind. However, CaVE+ completes its $10$ training epochs in roughly 2-3 minutes, whereas SPO+ and PFYL require more than 1 hour each, on average.
@@
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL
Deg $4$ 0.2004 0.0913 0.0913 0.0999 N/A
@@
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL
Deg $4$ 5.87 286.60 317.54 127.80 $\geq 100$h
This performance gap is even more pronounced for CVRP30 in Table 6(b). It takes roughly 20 seconds on average to solve a single CVRP30 instance, thus requiring 10 or so hours to traverse the entire dataset of 1,000 training instances once for SPO+ and PFYL. This makes end-to-end training with these methods impractical for real-world applications. In contrast, CaVE demonstrates its ability to handle such a challenging problem efficiently. Note that due to the scale of the problem, our experimental evaluation was not repeated with 10 random seeds and we used a smaller test set comprising only $10$ instances. All CaVE variants achieve test regrets of $9-10\%$ compared to the 2-Stage method’s far higher $20\%$, while requiring only 2-6 minutes of training each. To our knowledge, this is the hardest optimization problem ever targeted in the predict-then-optimize literature, a feat that is only possible due to the computational efficiency brought about by CaVE.

### 15 Mar 2024 v2

@@ 4.2 Cone-aligned Vector Estimation: three variants
$\mathcal{L}_{\text{CaVE}}(\hat{\bm{c}},\widetilde{\bm{A}}(\bm{c}))=-\text{cosine\_similarity}(\hat{\bm{c}},\bm{p}_{\hat{\bm{c}}})=-\dfrac{\hat{\bm{c}}^{\intercal}\bm{p}_{\hat{\bm{c}}}}{\lVert\hat{\bm{c}}\rVert\lVert\bm{p}_{\hat{\bm{c}}}\rVert},$ (4)
where, assuming there are $m$ binding constraints, i.e., $\widetilde{\bm{A}}(\bm{c})\in\mathbb{R}^{m\times d}$, the projection writes:
$\bm{p}_{\hat{\bm{c}}}={\widetilde{\bm{A}}(\bm{c})}^{\intercal}\bm{\lambda}^{*},\quad\bm{\lambda}^{*}=\operatorname*{arg\,min}_{\bm{\lambda}\geq\bm{0}}\lVert{\widetilde{\bm{A}}(\bm{c})}^{\intercal}\bm{\lambda}-\hat{\bm{c}}\rVert^{2}.$ (5)
@@
6:        Compute cosine similarity loss $\mathcal{L}_{\text{CaVE}}(\hat{\bm{c}},\widetilde{\bm{A}})$ (4)
@@
Figure 3 illustrates the three types of projection that can be employed in the CaVE framework. The exact projection projects the cost coefficients $\hat{\bm{c}}$ directly onto the surface of the optimal subcone; this is the approach that we have just laid out in Algorithm 1. The inner projection ensures that the projected vector lies strictly within the subcone. The heuristic projection is an approximation of the true projection onto the optimal subcone, used to reduce the computational cost. We will detail these three variants next.
@@
CaVE-E performs exact projection, wherein the optimal solution of the quadratic programming problem (5) is computed to locate the projection on the surface of the cone, as illustrated on the left in Fig. 3. Nevertheless, the CaVE-E method encounters a significant drawback due to its projection onto the face of the cone and the use of cosine similarity as the loss function. This approach results in the vanishing of gradients as the predicted cost vector nears the surface of the optimal subcone but is yet to enter it. Experimental evidence in Section 6 corroborates this issue – the regret of CaVE-E is typically higher than existing end-to-end methods, necessitating a modification.
@@
Due to the issue of vanishing gradients associated with CaVE-E, CaVE+ replaces the exact projection with what we refer to as “inner projection”. The goal is to obtain a projection of the predicted cost vector that lies inside the subcone. As all optimal projections will lie on a face of the subcone, we thus require a suboptimal solution to the projection problem. This is readily achieved by simply limiting the number of iterations in the quadratic programming solver and thus terminating prematurely. Since the solver uses the primal-dual interior point method, the feasibility is guaranteed at each iteration. In our experiments, the maximum number of iterations of the QP solver used by CaVE+ was set to $3$. A suboptimal projection will lie inside the optimal subcone, resulting in nonzero loss and a strong gradient signal that will push the ML model parameters to produce predictions that move toward the inside of the subcone. Compared to exact projection, this approach is more computationally efficient with fewer iterations of the interior point method. The inner projection is illustrated in the middle of Fig. 3.
@@
To alleviate the computational burden of repeated QP solving in both CaVE-E and CaVE+, a hybrid strategy is employed in CaVE-H. We interleave inner projections (obtained with a QP just as in CaVE+) with much cheaper heuristic projections. Unlike exact and inner projections, the heuristic projection does not necessitate solving a quadratic program and instead requires a simple convex combination of $\overline{\bm{A}}(\bm{c})$ and $\hat{\bm{c}}$ with weight $\gamma\in[0,1]$:
$\bm{p}_{\hat{\bm{c}}}=(1-\gamma)\hat{\bm{c}}+\gamma{\overline{\bm{A}}(\bm{c})}^{\intercal},$
where $\overline{\bm{A}}(\bm{c})=\frac{1}{m}\sum_{\bm{a}_{j}\in\widetilde{\bm{A}}(\bm{c})}{\bm{a}_{j}}$ is the average of all binding constraints. As illustrated on the right in Fig. 3, it is crucial to note that the heuristic projection is not guaranteed to be in the optimal subcone, but it still ensures that the cost coefficient vector is pushed in the direction of the optimal subcone. With probability $\beta<0.5$, i.e., in the minority of the iterations of Algorithm 1CaVE-H performs an inner projection via QP. With probability $(1-\beta)$, the heuristic projection is used instead, without any optimization required. In our experiments, $\gamma$ and $\beta$ are set to $0.2$ and $0.3$, respectively, and were not tuned any further for performance.
Method Iteration cost
PFYL $K\times$BLP
SPO+ 1×1$\times$1 ×BLP
NCE $\beta\times$BLP
CaVE-E 1×1$\times$1 ×QP
CaVE+ 1×1$\times$1 ×QP (partial)
CaVE-H $\beta\times$QP (partial)
Table 1: Comparison of state-of-the-art predict-then-optimize methods SPO+, PFYL, and NCE with the three CaVE variants w.r.t. “Iteration cost” (per training instance), i.e., the frequency of time-consuming solver calls required to compute the loss of a method during gradient descent. The methods are sorted in decreasing order of their iteration costs. SPO+ requires solving the BLP with predicted costs and PFYL requires solving the $K$ (typically 1-5) BLPs, whereas NCE with solution cache needs to solve only a small fraction ($100\times\beta\%$) of the BLPs. CaVE methods require solving a single QP, or partially solving a single QP in ($100\times\beta\%$) of the iterations.
@@
Table 1 summarizes the computational cost of training predict-then-optimize models using SPO+ [7], PFYL [3], NCE [20] and the three CaVE variants. We choose SPO+ and PFYL as they represent the state-of-the-art based on an extensive evaluation carried out in [28], though our experiments will include NCE with solution cache, a fast method, as well as accelerated variants of SPO+ and PFYL. The table shows that CaVE-H and NCE are the fastest to train, whereas PFYL and SPO+ exhibit similar costs, particularly when PFYL’s number of random perturbations of the predicted cost vector is set to $K=1$CaVE+ sits between CaVE-E and CaVE-H in terms of training cost. We will empirically examine these theoretical complexities in the next section.
@@ 5 Benchmark Datasets
The shortest path problem is a fundamental task in graph theory and network analysis, where the objective is to find a minimum weight path from a fixed source node to a fixed target node. Our instances are based on $5\times 5$ grid networks (SP5) where the source is the node in the northwest corner of the grid and the target is the node in the southeast corner. The cost coefficient $c_{ij}$ comes from
@@
$\Bigg{(}\frac{1}{\sqrt{10}}(\mathcal{B}\bm{x}_{i})_{j}+3\Bigg{)}^{\text{deg}}\cdot\epsilon_{ij},$
@@ 6.1 Experimental setup
The experiments were designed to evaluate both the training time and normalized regret, with the size of both the training and test sets being 1,000 instances. To account for randomness in data generation and stochastic model training, $5$ or $10$ random seeds are used to generate and train training/test sets and the corresponding ML models. Our comparative analysis encompassed several methods, including three variants of CaVE, a two-stage approach (least-squares regression on cost vectors), SPO+, PFYL, and NCE. A linear model was used as the cost prediction model class for all the aforementioned approaches; this is a standard choice in the literature starting with [7]. In the case of PFYL, the size of random sampling is fixed at $K=1$; for NCE, the solving ratio is set at $\beta=5\%$.
The Adam optimizer was used for gradient descent for all our experiments. For SP5, a learning rate of $0.01$ was used with all methods for 10 epochs, except for the two-stage and NCE method which allotted 20 epochs at the same learning rate. For both TSP and CVRP, the only modification was to increase the learning rate to $0.05$, maintaining the number of epochs at 10. In selecting the number of epochs, careful consideration was given to the convergence of each model, ensuring that our results are not skewed by over- or under-training. The loss curves in Fig. 4 show that all methods exhibit a convergent behavior, with CaVE methods doing so earlier than the baselines. Note that the 2-Stage models are trained essentially to its optimality as they are convex optimization problems.
Figure 4: Validation regeret curves for TSP50 with a polynomial of degree 4 (left) and 6 (right). The vertical axis represents the average normalized regret values of each of the five methods for the validation dataset with $1000$ instances.
@@ 6.2 Results
Tables 2(b)3(b)4(b)5(b), and 6 summarize the results. The tables correspond to the Shortest Path problem on $5\times 5$ grid (SP5), TSP with 20 nodes (TSP20), TSP with 50 nodes (TSP50), CVRP with 20 customers (CVRP20), and CVRP with 30 customers (CVRP30), respectively. Each table contains two sub-tables: (a) reports the normalized regret metric defined in Section 3.2 on 1,000 test instances and (b) reports the training time in seconds; both quantities are averages over $5$ random seeds for CVRP20 and $10$ random seeds for others with standard deviation, while the experiments for CVRP30 is not repeated due to running time. Each sub-table has two rows, one corresponding to datasets that use a degree-4 polynomial in the (unknown) mapping from instance features to cost coefficients, and the other for a degree-6 polynomial; we refer to Section 5 for details on the role of the polynomial in the function that we are attempting to learn, but note that the higher the degree the more difficult the learning task. In each sub-table, we bold the best-performing method, excluding the 2-Stage method as it is often fast to train, but has much worse regret than end-to-end methods.
@@
This is the easiest BLP (in fact, LP) we will look at. It serves as a sanity check for any new method in this space. Table 2(b) shows that SPO+ and PFYL achieve the lowest test regrets, closely followed by CaVE-H and CaVE+. The latter two trains roughly $10$ times faster than SPO+ and PFYL, already substantiating our claim that QP solving is faster in solving the optimization problem itself, even for small-scale polynomial-time solvable shortest path problems. In addition, although NCE is fast, it has a higher regret than others.
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
@@
Deg $4$ $8.82\pm 1.15$ $10.73\pm 1.54$ $8.39\pm 0.95$ $8.35\pm 0.88$ $7.79\pm 1.00$ $\bm{7.68}\pm 0.99$ $11.34\pm 1.11$
Deg $6$ $12.58\pm 2.14$ $11.30\pm 1.30$ $8.89\pm 0.90$ $8.84\pm 1.00$ $\bm{7.72}\pm 1.11$ $7.86\pm 0.96$ $13.78\pm 1.58$
(b) Average Training Time (Sec) with Standard Deviation
@@
For the TSP, SPO+, PFYL and NCE employ the Dantzig–Fulkerson–Johnson (DFJ) formulation [5] to solve this BLP efficiently. For both TSP20 and TSP50, CaVE+ achieves the best time-regret trade-off across all methods: its regret is the best or second-best across all methods and its training time is the second-best after CaVE-H. As shown in part (b) of Tables 3(b) and 4(b)CaVE+ trains in roughly half the time of SPO+ and PFYL, achieving the same or even better test regret. Additionally, we also compared our methods with SPO+ Rel and PFYL Rel, which employ a linear relaxation of the BLP during training, as detailed in the Appendix 0.A.1. While they do reduce the training time of vanilla SPO+PFYL by a bit, this typically comes at an increase in regret. Similarly, NCE has a prohibitively high regret.
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
@@
Deg $4$ $12.12\pm 0.89$ $7.35\pm 0.40$ $6.20\pm 0.24$ $7.69\pm 0.33$ $\bm{5.95}\pm 0.16$ $6.56\pm 0.21$ $12.21\pm 0.88$
Deg $6$ $21.32\pm 1.81$ $8.01\pm 0.45$ $\bm{6.97}\pm 0.37$ $9.52\pm 0.64$ $7.48\pm 0.36$ $7.41\pm 0.37$ $14.31\pm 0.40$
(b) Average Training Time (Sec) with Standard Deviation
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
@@
Deg $4$ $28.16\pm 1.08$ $15.19\pm 0.65$ $7.69\pm 0.22$ $9.59\pm 0.44$ $\bm{7.57}\pm 0.20$ $8.03\pm 0.23$ $14.31\pm 0.40$
Deg $6$ $52.61\pm 2.36$ $23.25\pm 2.41$ $\bm{8.57}\pm 0.38$ $11.28\pm 0.80$ $10.26\pm 0.46$ $9.00\pm 0.52$ $17.12\pm 0.48$
(b) Average Training Time (Sec) with Standard Deviation
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL NCE
Deg $4$ $10.10\pm 0.64$ $9.26\pm 1.56$ $6.44\pm 0.24$ $7.92\pm 0.52$ $\bm{5.94}\pm 0.25$ $6.32\pm 0.28$ $15.77\pm 0.96$
Deg $6$ $19.50\pm 1.22$ $11.64\pm 0.25$ $\bm{7.94}\pm 0.54$ $11.44\pm 1.14$ $8.75\pm 0.28$ $8.09\pm 0.57$ $18.96\pm 1.01$
(b) Average Training Time (Sec) with Standard Deviation
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL NCE
Deg $4$ $1.65\pm 0.48$ $213.56\pm 42.36$ $153.56\pm 11.08$ $\bm{44.52}\pm 6.27$ $7020.11\pm 1043.05$ $3773.31\pm 288.84$ $583.56\pm 170.67$
Deg $6$ $1.54\pm 0.25$ $208.95\pm 12.90$ $127.94\pm 13.84$ $\bm{51.83}\pm 8.78$ $2204.83\pm 99.86$ $6197.84\pm 288.63$ $470.20\pm 84.46$
As mentioned earlier, we are the first to tackle a CVRP in an end-to-end predict-then-optimize setting. To solve CVRP as a BLP, we formulated the problem with k-path cuts [12] and solved it using Gurobi. For CVRP20, CaVE+, SPO+ and PFYL compete for the lowest regret, with CaVE-H running closely behind. However, CaVE+ completes its $10$ training epochs in roughly 2-3 minutes, whereas SPO+ and PFYL require more than 1 hour each, on average. The regret of NCE is high and the training time has increased a lot. Similar as TSP, we also employ a linear relaxation of the BLP during training for SPO+ and PFYL (see Appendix 0.A.2), which results in a significant decrease in training time at the cost of regret.
@@
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL NCE
Deg $4$ 0.1972 0.1254 0.0913 0.0999 N/A 0.1828
@@
Methods 2-Stage CaVE-E CaVE+ CaVE-H SPO+ PFYL NCE
Deg $4$ 9.27 331.73 287.77 132.62 $\geq 100$h 884.95
This performance gap is even more pronounced for CVRP30 in Table 6. It takes roughly 20 seconds on average to solve a single CVRP30 instance, thus requiring 8 or so hours to traverse the entire dataset of 1,000 training instances once for SPO+ and PFYL. This makes end-to-end training with these methods impractical for real-world applications. In contrast, CaVE demonstrates its ability to handle such a challenging problem efficiently. Note that due to the scale of the problem, our experimental evaluation was not repeated with random seeds and we used a smaller test set comprising only $10$ instances. All CaVE variants achieve test regrets of $9$-$12\%$ compared to the 2-Stage method’s far higher $20\%$, while requiring only 2-6 minutes of training each. In addition, NCE achieves $18\%$ with $15$ minutes of training and is thus worse than all variants  CaVE in both metrics. To our knowledge, this is the hardest optimization problem ever targeted in the predict-then-optimize literature, a feat that is only possible due to the computational efficiency brought about by CaVE.